Circular No. 8922

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2006 SF_{369}

K. S. Noll and S. D. Kern, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; H. F. Levison, Southwest Research Institute; and E. A. Barker, STScI, report that the 1:3 Neptune-librating transneptunian object 2006 SF_{369} (cf. MPEC 2006-U61) is a binary. One 260-s exposure was made at each of four dithered positions on the detector during 2007 Nov. 13.3563–13.3868 UT with the Planetary Camera of the Wide Field Planetary Camera 2 on the Hubble Space Telescope (HST), using the F606W filter (wide V); these exposures show two components for 2006 SF_{369} that were separated by an angular distance of 0′′.109 ± 0′′.003 and were nearly equal in brightness (differing by < 0.1 mag). The fainter component was located at −0′′.019 ± 0′′.002 in α and −0′′.107 ± 0′′.002 in δ from the brighter component. One of three objects known or suspected to be in 1:3 resonance with Neptune [the others being (136120) 2003 LG_{7} (MPEC 2005-L33) and 2005 EO_{297} (Deep Ecliptic Survey integration)] and which have been observed with HST, 2006 SF_{369} is the only detected binary.

(119067) 2001 KP_{76} AND (160091) 2000 OL_{67}

F. Marchis and M. Baek, Carl Sagan Center, SETI Institute; and J. Berthier, P. Descamps, and F. Vachier, Institut de Mécanique Céleste et de Calcul des Éphémérides, Paris; jointly with K. S. Noll and S. D. Kern, Space Telescope Science Institute; and W. M. Grundy, Lowell Observatory, report on companions to the cubewanos (119067) 2001 KP_{76} [cf. MPEC 2001-M60, MPO 88168] and (160091) 2000 OL_{67} [cf. MPEC 2000-T41, MPO 121405]. Observations of (119067) were made during 2007 May 8.576–8.597 UT with the HST as noted above; the processed image (the result of four dithered positions on the detector, as described above) shows two components separated by 0′′.29 in p.a. 276^{o} and having a brightness difference of only 0.1 magnitude. A companion was detected around (160091) in HST observations taken as above between 2007 June 26.514 and 26.487; a brightness difference of 0.6 magnitude is clearly visible for the two components in the processed image (processed as above), and its separation is 0′′.26 in p.a. 277^{o}, corresponding to a projected distance of 7800 km.

COMET C/2007 T1 (McNAUGHT) Total visual magnitude estimates by D. A. J. Seargent, Cowra, N.S.W.

(25×100 binoculars): 2007 Nov. 10.42 UT, 9.7; 2008 Jan. 8.45, 8.7; 11.48, 8.3; Feb. 7.48, 8.7; 9.47, 8.8.

2008 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green